OpenStudy (anonymous):

(x+8)(x-13)(x+5)>0 {xI_} if there is a solution is the solution all real numbers or is there no solution?

OpenStudy (anonymous):

jst use the number line to find the solution solve for x (x+8)(x-13)(x+5)=0 x=-8, x=13, x=-5 and then choose test points

OpenStudy (anonymous):

So there wold be no solution?

hero (hero):

There is a solution, but it is limited to an interval. It isn't all real numbers though

myininaya (myininaya):

---|----|----|---- -8 -5 13 choose a number before -8 like -9 and plug it into (x+8)(x-13)(x+5) choose a number between -8 and -5 like -6 and plug it into (x+8)(x-13)(x+5) choose a number between -5 and 13 like 0 and plug it into (x+8)(x-13)(x+5) choose a number after 13 like 14 and plug it into (x+8)(x-13)(x+5)

myininaya (myininaya):

(-9+8)(-9-13)(-9+5)=(-1)(-22)(-4)=negative number<0 (-6+8)(-6-13)(-6+5)=(2)(-19)(-1)=positive number>0 (0+8)(0-13)(0+5)=(8)(-13)(5)=negative number<0 (14+8)(14-13)(14+5)=(22)(1)(19)=positive number>0

myininaya (myininaya):

so which intervals >0 ?